# Magnetic field dynamics in presence of Hall conductivity, and thermodiffusion

### v1

@@ 2 Equations in the presence of thermodiffusion, and external electro-moving force (EMF)
$\frac{1}{c}\frac{{\partial\bf B}_{in}}{\partial t}=-{\bf\nabla\,\times\,E},\quad{\bf\nabla\cdot B}_{in}=0,\quad{\bf j_{in}}=\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{in},\quad\qquad\qquad$ (28)
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]},\qquad$ (29)
${\bf j}_{in}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr{]},\quad{\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]},\quad$ (30)
$\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{ex}^{v}={\bf j_{ex}},\quad{\bf\nabla\cdot B}_{ex}=0,\qquad{\bf B}_{ex}={\bf B}_{ex}^{v}+{\bf B}_{ex}^{vf},\qquad\qquad$ (31)
${\bf j}_{ex}=\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr{]}\qquad\qquad\qquad$ (32)
$+\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl{[}{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}]}\biggr{]},\qquad\qquad\qquad$
@@
Here ${\bf E}$, ${\bf E^{\prime}}$ are inner electric fields measured in the laboratory, and comoving coordinate system, respectively. In addition there are given external electrical field ${\bf E^{{}^{\prime\prime}}}$, external current ${\bf j_{ex}}$, creating a vortex external magnetic field ${\bf{B_{ex}^{v}}}$; and independent vortex-free magnetic field ${\bf{B_{ex}^{vf}}}$.
@@
There is a complicate situation due to presence of a external fields, currents and thermodiffusion. All currents are producing magnetic fields, which, together with the external field ${\bf{B_{ex}^{vf}}}$, are taking part in the formation of Hall currents. The currents are produced in different processes, but the magnetic field which they have created, is acting together, as one vector ${\bf B}$. These properties are represented in the system of equations (28)-(33).
3 Simple models with an action of the Hall current in the presence of thermodiffusion
@@ 3.1 Magnetized plasma cylinder
Here we obtain from (28) - (33) the equations, describing the behavior of the plasma cylinder (see Fig. 1) with a uniform magnetic field $B_{0}$ along its axis, and a radial temperature gradient. In the case of cylinder symmetry $\frac{\partial}{\partial z}=\frac{\partial}{\partial\phi}$, the only non-zero parameters of heat flux and electric current left are: $q_{r},q_{\phi},j_{r},j_{\phi}$.
@@
In absence of the external electric field $E^{{}^{\prime\prime}}$ and mutually perpendicular vectors ${\bf B,\nabla}T$, equation (32) is written as follows:
${\bf j_{ex}}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}({{\bf\nabla}T}-\frac{\omega\tau_{e}}{B}[{\bf B\times\nabla}T]).$ (34)
The components of the electrical current density $j_{ex}$ in a cylinder with $B_{z}$ and radial temperature gradient are written as:
$j_{ex,r}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr},\qquad j_{ex,\phi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (35)
From eq. (31) we have the equation for the inner magnetic field, induced by the Hall current, as:
$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\lambda_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{dr}.$ (36)
Solution of this equation at different input parameters was obtained in [21]. The magnetic field $B_{z}^{v}$, induced by the Hall current due to thermodiffusion, in all cases had a sign, opposite to the external field $B_{z,ex}^{v}$.
When the external radial electric field $E_{r,ex}$ is present in the cylinder, instead of the radial temperature gradient, then instead of the Eq.(36), we have the following equation for the induced field $B_{z}^{v}$:
$\quad\frac{dB_{z}^{v}}{dr}=\frac{4\pi}{c}\frac{\sigma_{0}\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}E_{r,ex},\quad\sigma_{0}=\frac{32}{3\pi}\frac{e^{2}n_{e}\tau_{e}}{m_{e}}.$ (37)
Here $\sigma_{e}$ is the scalar plasma electrical conductivity. This field has the same opposite sign in relation to the external field $B_{z,ex}^{vf}$.
@@ 3.2 Magnetized plasma torus
Figure 2: Torus with a initial electric field $E_{0}$, that produce circular magnetic field $B_{0}$. $B_{0}$ and temperature gradient $\nabla T$ create Hall electric current $j_{Hall}$ in opposite direction to the $E_{0}$. The induced magnetic field $B_{1}$ is determined by the Hall current $j_{Hall}$.
Let us consider a circular plasma torus, in the center of which there is a linear electrical current with a linear density ${J_{z}}$, directed perpendicular to the symmetry plain of the torus $(x,y)$, see Fig.2. The cross section of the torus, which includes $z$-axis, is supposed to be circular, with a radius $R_{t}$. In the symmetry plane $z=0$ the torus is situated between the radius $R_{out}$ and $R_{in}$, where:
@@
Torus surface is defined by the 4-th order equation. It is formed by a rotation of a circle with the radius $R_{t}$, around a $z$-axis, when the center of the circle has the radius $R=R_{in}+R_{t}$. The equation of this surface is written as follows
$(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=R_{t}^{2},\quad{\rm which\,\,reduced\,\,to}\quad$
$(x^{2}+y^{2}+z^{2}+R^{2}-R_{t}^{2})^{2}=4R^{2}(x^{2}+y^{2}).\,\,\,$ (39)
@@
By heating the central ring of the torus and formation of the temperature gradient ${\bf\nabla T}$, we produce a heat flux inside the torus directed to the surface of the torus in the plane, containing the $z$ axis. In the cylinder coordinate system $r,\phi,z$ the non-zero temperature gradient takes place along the coordinate $\rho$, defined as:
$\rho^{2}=(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=(r-R)^{2}+z^{2},\quad 0\leq\rho\leq R_{t}.$ (40)
With this definition we have  $\rho d\rho=(r-R)dr+zdz$,
$\frac{\partial T}{\partial r}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial r}=\frac{r-R}{\rho}\frac{dT}{d\rho},\qquad\frac{\partial T}{\partial z}=\frac{dT}{d\rho}\frac{\partial\rho}{\partial z}=\frac{z}{\rho}\frac{dT}{d\rho}.$ (41)
@@
For constant heat flux $Q$ through the unit length of the torus, the temperature gradient along the small torus radius $\rho$ is determined as
$q=\frac{Q}{2\pi}=\frac{\kappa_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\quad\kappa_{0}=\frac{320}{3\pi}\frac{k^{2}Tn_{e}\tau_{e}}{m_{e}}.$ (42)
@@
Here the scalar heat conductivity coefficient $\kappa_{0}$ is determined in [17]. The thermodiffusion is forming an electrical current along the temperature gradient over the $\rho$ coordinate $j_{ex,\rho}$, and a hall current $j_{ex,\psi}$ along the small circle of the torus:
$j_{ex,\rho}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\qquad j_{ex,\psi}=-\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho}.$ (43)
@@
Due to the axial symmetry, the magnetic field in the torus is produced only by its Hall component $j_{ex,\psi}$, namely, by the circular current along the small circle of the torus, see Fig.2. The torus with such current is used in the construction of the devices for obtaining controlled thermonuclear reaction [23], where such currents are concentrated in wire around the torus, creating the toroidal field going around the torus. In our model similar currents of Hall origin are created in the plasma of the torus itself, when the current is created in the wire circles, with the current $I$, along the torus, with $n$ wire circles on the unit length. The total circular current around the small torus radius is equal to $2\pi RnI$. The magnetic field circulation along a big circle on the distance $r$ from the central big circle is equal to $2\pi rB$. From the circulation theorem we obtain then the poloidal field inside such torus as [24]:
$B_{\phi}(r)=nI\frac{R}{rc},\quad R_{in}<r<R.$ (44)
@@
In the plasma torus there is a circular Hall current density $j_{ex,\psi}$, defined in (43). The magnetic field formed by the Hall current could be considered as sum of thin circular currents, similar to the wire current in the laboratory torus, considered above. The total circular current on the radial distance $r_{0}$ from the center is equal to ($j_{ex,\psi}2\pi r_{0}d\rho$). The magnetic field circulation from this current along the torus at the radius $r<r_{0}$ is equal to $2\pi rB$. So the magnetic field $dB_{\phi}$ produced by this sheet of current at the radius $r$ is written as:
$dB_{\phi}(r)=j_{ex,\psi}\frac{r_{0}}{rc}d\rho,\quad R_{in}<r<r_{0}.$ (45)
@@
It follows from (44), after substituting ($nI$) by ($j_{ex,\psi}(\rho)d\rho$), and $R$ by $r_{0}$. The Hall magnetic field from the whole current in the plasma torus is obtained by integration of (45) over the torus thickness as
$B_{\phi}(r)=\int_{0}^{R-R_{in}}j_{ex,\psi}\frac{r_{0}}{rc}d\rho,=\frac{1}{rc}\int_{0}^{R-R_{in}}j_{ex,\psi}r_{0}d\rho,\qquad R_{in}<r<R.$ (46)
@@
On the symmetry plane $z=0$ there is a connection $\rho=R-r_{0}$, $d\rho=-dr_{0}$. For constant current density we have after integration:
[…]

### v2

@@ 3.2 The system of MHD equations, related to magnetic field behavior
$\frac{1}{c}\frac{{\partial\bf B}_{in}}{\partial t}=-{\bf\nabla\,\times\,E},\quad{\bf\nabla\cdot B}_{in}=0,\quad{\bf j_{in}}=\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{in},\qquad$ (34)
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}\qquad$
$-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}-\frac{{\bf\nabla}P_{e}}{en_{e}},\quad\qquad$ (35)
${\bf j}_{in}=\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{\prime}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{\prime})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{\prime}]}\biggr],\qquad$ (36)
${\bf E^{\prime}}={\bf E}+\frac{1}{c}{{\bf[v\times B]}}+\frac{{\nabla}P_{e}}{en_{e}},\,\,\quad\qquad$ (37)
$\frac{c}{4\pi}{\bf\nabla\,\times\,B}_{ex}^{v}={\bf j_{ex}},\quad{\bf\nabla\cdot B}_{ex}=0,\quad{\bf B}_{ex}={\bf B}_{ex}^{v}+{\bf B}_{ex}^{vf},\qquad$ (38)
${\bf j}_{ex}=\frac{\lambda_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf\nabla}T+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot\nabla}T){\bf B}-\frac{\omega\tau_{e}}{B}{\bf[B\times\nabla}T]\biggr]\quad$ (39)
$+\frac{\sigma_{0}}{1+\omega^{2}{\tau_{e}}^{2}}\biggl[{\bf E^{{}^{\prime\prime}}}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}{\bf(B\cdot E^{{}^{\prime\prime}})B}-\frac{\omega\tau_{e}}{B}\bf{[B\times E^{{}^{\prime\prime}}]}\biggr],\hskip 22.99988pt$
@@
Here $\bf E$}, $\bf E^{\prime}$} are inner electric fields measured in the laboratory, and in the co-moving coordinate system, respectively. In addition there are given external electrical field $\bf E^{{}^{\prime\prime}}$}, external current $\bf j_{ex}$}, creating a vortex external magnetic field $\bf{B_{ex}^{v}}$}; and independent vortex-free magnetic field $\bf{B_{ex}^{vf}}$}.
@@
There is a complicate situation due to presence of a external fields, currents and thermo-diffusion. All currents are producing magnetic fields, which, together with the external field $\bf{B_{ex}^{vf}}$}, are taking part in the formation of Hall currents. The currents are produced in different processes, but the magnetic field which they have created, is acting together, as one vector $\bf B$}. These properties are represented in the system of equations (34)-(40).
4 Several applications
4.1 Thrusters
Thrusters are reactive engines at low power, and very high efficiency, based on the electromagnetic acceleration of plasma. A conception of this engine was suggested in the year 1957 by A.I. Morozov [22]. A simplified version of their action is based on the consideration of the ideal plasma behavior inside the mutually perpendicular external electric $\bf E^{{}^{\prime\prime}}$} and magnetic $\bf B_{ex}$} fields, where the internal electrical field in the laboratory system $\bf E$} is zero. For the ideal plasma with infinite conductivity $\sigma_{0}$ we have from (3.2) the relation, determining the equilibrium plasma velocity v in these conditions
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}=0.$ (41)
This equation has an exact solution
${\bf v}=c\frac{\bf E^{{}^{\prime\prime}}\times B}{B^{2}}.$ (42)
As followed from (34), at zero $\bf E$} the inner magnetic field $\bf B_{in}$}=0 therefore, we take here ${\bf B}={\bf B_{ex}}$.
The review’s devoted to the 30th anniversary of small plasma thrusters (SPT) operation in space are given in Refs. [23] (about the conceptual development of stationary plasma thrusters), and [24] (about stationary plasma thrusters operate in space).
Usually we have situations when $E\ll B$. so that the speed of the out flowing plasma $v\ll c$. If, by some reason, the value of $\bf E$} is approaching that of $\bf B$}, the speed of the out flowing plasma is approaching light speed $c$. In this situation the approximate Lorentz transformation formulae should be substituted by its exact variant [25], and instead of (41), (42) we obtain following relations for the geometry where, in Cartesian coordinates $(x,y,z)$
${\bf v}=(V,0,0),\quad{\bf E}=(0,E,0),\quad{\bf B}=(0,0,B),$ (43)
the following result
$E=-\frac{V}{c}\frac{B}{\sqrt{1-V^{2}/c^{2}}}-E^{{}^{\prime\prime}}=0.$ (44)
This equation has an exact solution
$V={c}\frac{E^{{}^{\prime\prime}}}{\sqrt{B_{ex}^{2}+{E^{{}^{\prime\prime}}}^{2}}}.$ (45)
4.2 Neutron star crust
Biermann battery effects connected with a magnetic field behavior, and its possible role in the seed magnetic field creation in the neutron star (NS) crust, have been considered in [26]. The authors used transport coefficients from the works [27], [28],[29], where numerical estimations of some integral have been made. Here we obtain a fully analytic solution of this problem, using some approximation.
In the middle aged NS crust a temperature is moderate, and a density is very high, so the electrons are strongly degenerate and ultrarelativistic. Following [27], we extend approximately the application of the Eq. (28) to the relativistic electrons, by using in the transport coefficients, the effective electron mass $m_{*}$, instead of $m_{e}$, with
$m_{*}\approx{m_{e}}{\sqrt{1+\frac{p_{fe}^{2}}{(m_{e}c)^{2}}}},\qquad p_{fe}=\bigg(\frac{3n_{e}}{8\pi}\biggr)^{1/3}h.$ (46)
In this case the coefficients of conductivity and thermo-diffusion, with account of (46), are taken in the form
$\sigma_{0}^{du}=\frac{e^{2}n_{e}}{m_{*}}\tau_{du},\qquad\lambda_{0}^{du}=\frac{4\pi^{3}e}{(9\pi)^{1/3}}\frac{n_{e}^{1/3}k^{2}T}{h^{2}}\tau_{du},$ (47)
with
$\tau_{du}=\frac{3}{32\pi^{2}}\frac{h^{3}n_{e}}{Z^{2}e^{4}m_{*}n_{N}\Lambda},\qquad h\,\,\,\,{\mbox{ is the Planck constant}}.$ (48)
The ratio of transport coefficients is written as
$\frac{\lambda_{0}^{du}}{\sigma_{0}^{du}}=\frac{k}{e}\,\frac{4\pi^{8/3}}{3^{2/3}}\frac{kTm_{*}}{h^{2}n_{e}^{2/3}},$ (49)
with possible non-zero value of $\nabla n_{e}\times\nabla T\neq 0.$
For the baro-diffusion term in (7), (3) of degenerate electrons the equation of state in the relativistic limit is written in the form (see [16])
$P_{d}=\frac{1}{8}\biggl(\frac{3}{\pi}\biggr)^{1/3}c\,h\,n_{e}^{4/3}\biggl[1+\frac{2\pi^{2/3}}{3^{5/3}}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}n_{e}^{-2/3}\biggr]$ (50)
The creation of the seed magnetic field in the strongly degenerate plasma is described by the equation (28) We obtain from (49) and (50)
$\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)=\frac{k}{e}\,\frac{4\pi^{8/3}}{3^{2/3}}\frac{km_{*}}{h^{2}}\biggl(\frac{\nabla T}{n_{e}^{2/3}}-\frac{2}{3}\frac{T\nabla n_{e}}{n_{e}^{5/3}}\biggr),$ (51)
$\nabla\times\biggl(\frac{\nabla P_{d}}{n_{e}}\biggr)=-\frac{\nabla n_{e}}{n_{e}^{2}}\times{\nabla P_{d}}=-\frac{\pi^{1/3}ch}{2\cdot 3^{4/3}n_{e}^{4/3}T}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}(\nabla n_{e}\times{\nabla T}).$ (52)
With account of (51),(52) we have from (28)
$\frac{\partial{\bf B}}{\partial t}=-\biggl[\frac{ck}{e}\,\frac{8\pi^{8/3}}{3^{5/3}}\frac{kTm_{*}}{h^{2}n_{e}^{5/3}}+\frac{c}{e}\,\frac{\pi^{1/3}ch}{2\cdot 3^{4/3}n_{e}^{4/3}T}\biggl(\frac{2\pi kT}{h\,c}\biggr)^{2}\biggr](\nabla n_{e}\times{\nabla T})$ (53)
This equation is rewritten as (see also [26])
$\frac{\partial{\bf B}}{\partial t}=-\frac{ck}{e}\,\frac{2\pi^{7/3}}{3^{4/3}n_{e}^{4/3}}\biggl(\frac{kT}{h\,c}\biggr)\biggl(1+\frac{4\pi^{1/3}}{3^{1/3}}\frac{m_{*}c}{hn_{e}^{1/3}}\biggr)(\nabla n_{e}\times{\nabla T}).$ (54)
In this consideration the baro-diffusion term is based on exact equation of state for relativistic electrons, while in the thermo-diffusion term we used corrected non-relativistic kinetic coefficients, similar to [27]. In this situation the physical interpretation of terms in the brackets of (54) is not so clear.
@@ 4.3 Magnetized plasma cylinder
The model of the magnetized plasmic cylinder with a radial temperature gradient was investigated in [20]. The temperature gradient created the thermo-diffusion electrical current, in the external uniform magnetic field along the cylinder axis (see Fig.1). The equations describing the structure of the magnetic field in presence of the external thermo-diffusion force, used in [20], follow from Eqs. (34) - (40).
It was obtained in [20], that thermo-diffusional Hall currents in this situation create magnetic field of the sign, opposite the the external one, and can substantially decrease the resulting magnetic field. This behavior is in accordance with the Lenz’s law [30] in electrodynamic.
@@ 4.4 Magnetized plasma torus
The plasma torus configurations are used in construction of devices for thermonuclear energy production (Tokamak, Stellarator) [31],[32]. Bilding of ITER (International Thermonuclear Experimental Reactor) is based on the Tokamak intallation [33].
In astrophysics the plasma torus may be formed around compact objects (white dwarfs, neutron stars, black holes) during accretion in the binary system. The giant magnetized plasma torus configuration, formed around the central nuclei in the active galactic nuclei and quasars, are considered now as a most plausible model, explaining many different observational features [34],[35].
Here we consider the structure of a circular plasma torus, using Eqs. (34) -(40). Through the center of the torus there is a linear electrical current, situated perpendicular to the symmetry plain of the torus $(x,y)$, with a linear current $J_{z}^{ex}$}, see Fig.2. The cross section of the torus, perpendicular to $z$-axis, is supposed to be circular. In the symmetry plane $z=0$ the torus occupies the space between the radius $R_{out}$ and $R_{in}$. The cross section of the torus along $z$-axis is also taken as a circular, with the radius $R_{t}$, so that
@@
Figure 2: Torus with a initial electric current $j_{0}$, that produce circular magnetic field $B_{0}$. $B_{0}$ and temperature gradient $\nabla T$ create Hall electric current $j_{Hall}$. The induced magnetic field $B_{1}$ is determined by the Hall current $j_{Hall}$ and is oriented in the opposite direction to the $B_{0}$.
@@
The torus surface is defined by the 4-th order equation. It is formed by a rotation of a circle with the radius $R_{t}$, around a $z$-axis, when the center of the circle has the radius $R=R_{in}+R_{t}$. The equation of this surface is written as follows
$(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=R_{t}^{2},\quad{\rm which\,\,is\,\,reduced\,\,to}\quad$
$(x^{2}+y^{2}+z^{2}+R^{2}-R_{t}^{2})^{2}=4R^{2}(x^{2}+y^{2}).\,\,\,\quad$ (56)
@@
By heating the central ring of the torus and formation of the temperature gradient $\nabla T$}, a heat flux is created inside the torus along its smaller radius $\rho$, directed to the surface of the torus. In the cylinder coordinate system $r,\phi,z$ the the coordinate $\rho$ is defined as:
$\rho^{2}=(\sqrt{x^{2}+y^{2}}-R)^{2}+z^{2}=(r-R)^{2}+z^{2},\quad 0\leq\rho\leq R_{t},$ (57)
@@
so that in the plane $z=0$,  $\rho d\rho=(r-R)dr$, for $R_{in}<r<R$. For a constant heat flux $Q$ through the unit length of the large circle of the torus, the temperature gradient along the small torus radius $\rho$ is determined as
$q=\frac{Q}{2\pi\rho}=-\frac{\kappa_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\quad\kappa_{0}=\frac{320}{3\pi}\frac{k^{2}Tn_{e}\tau_{e}}{m_{e}}.$ (58)
@@
Here the scalar heat conductivity coefficient $\kappa_{0}$ is determined in [16]. The thermo-diffusion is forming the electrical current along the temperature gradient along the $\rho$ coordinate $j_{in,\rho}$, and a hall current $j_{in,\psi}$ along the small circle of the torus:
$j_{in,\rho}=\frac{\lambda_{0}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho},\qquad j_{in,\psi}=\lambda_{0}\frac{\omega\tau_{e}}{1+(\omega\tau_{e})^{2}}\frac{dT}{d\rho}.$ (59)
Due to the axial symmetry, the magnetic field inside the torus is produced only by the Hall component $j_{in,\psi}$,corresponding to a circular current along the small circle of the torus, see Fig.2.
@@
The torus with such current is used in the construction of the devices for obtaining controlled thermonuclear reaction [32], where these currents are concentrated in the wire wound up in the torus around its small radius. Such current is creating the toroidal field along the big circle the torus. In our model the external toroidal magnetic field is created by the central linear current.
To calculate the magnetic field created by the thermo-diffusion $\psi$ - Hall current component, we use the way similar to the one, which used for calculation of toroidal field created by the current through the wound up wire. If $I$ is the current along the wire, and $n$ is the number of wire small circles on the unit length of the torus big circle, than the external circular magnetic field $B_{\varphi}(r)$ inside the torus, at $R_{t}\ll R$, is approximately written as [36]
$B_{\varphi}^{in}(r)=\frac{4\pi}{c}nI\frac{r}{R},\quad r\approx R.$ (60)
@@
In the plasma torus there is a circular Hall current density $j_{in,\psi}$, defined in (59). The magnetic field formed by the Hall current could be considered as sum of thin circular currents, similar to the wire current in the laboratory torus, considered above. The $\psi$ - current in the torus on the radial distance $r$ from the center is equal to ($j_{in,\psi}2\pi rd\rho$). The magnetic field circulation produced by this current along the large circle of the torus torus, is approximately equal to $2\pi R\,dB_{\varphi}^{in}(r)$. So the magnetic field $dB_{\phi}^{in}$ produced by this sheet of current at the radius $r$ is written as:
$dB_{\varphi}(r)=-\frac{4\pi}{c}j_{in,\psi}\frac{r}{R}d\rho,\quad R_{in}<r<R.$ (61)
@@
It follows from (60), after substituting ($nI$) by ($j_{ex,\psi}(\rho)d\rho$). The Hall magnetic field from the whole current in the plasma torus is obtained by integration of (61) over the torus thickness, as
$B_{\varphi}(r)=-\frac{4\pi}{c}\int_{R-r}^{R-R_{in}}j_{in,\psi}\frac{r}{R}d\rho,=-\frac{4\pi}{Rc}\int_{R-r}^{R-R_{in}}j_{in,\psi}rd\rho,\qquad R_{in}<r<R.$ (62)
@@
On the symmetry plane $z=0$ there is a connection $\rho=R-r$, $d\rho=-dr$. For a constant current density we have after integration:
$B_{\varphi}^{in}(r)=-\frac{4\pi}{Rc}\int_{R-r}^{R-R_{in}}j_{in,\psi}(R-\rho)d\rho=-\frac{2\pi}{Rc}j_{in,\psi}(r^{2}-R_{in}^{2}).\quad R_{in}<r<R.$ (63)